# Giulia Gaggero and Elisa Gorla

### v2

@@ 3 Random polynomial systems
For any system $\mathcal{F}$ of equations of degree at least $q$, one has
$\mathcal{F}\cup\{x_{1}^{q}-x_{1},\ldots,x_{n}^{q}-x_{n}\}\in\mathcal{U}_{\{q,\ldots,q\}}.$
Unlike semiregular systems [23], algebraically random systems exist for any choice of the parameters and over every finite field.
Remark 14.
Algebraic random systems over $\mathbb{F}_{q}$ exist for any choice of $n,m$ and $1\leq d_{1}\leq\ldots\leq d_{m}$. This corresponds to the existence of regular sequences of any given degrees in $\mathbb{F}_{q}[x_{1},\ldots,x_{n}]$, as any system of $m$ equations in $\mathbb{F}_{q}[x_{1},\ldots,x_{n}]$ of degrees $d_{1},\ldots,d_{m}$ which contains a regular sequence in degrees $d_{1},\ldots,d_{n}$ is algebraically random. Some regular sequences in $\mathbb{F}_{q}[x_{1},\ldots,x_{n}]$ of degrees $d_{1},\ldots,d_{n}$ are for example
$x_{1}^{d_{1}}+g_{1},x_{2}^{d_{2}}+g_{2},\ldots,x_{n}^{d_{n}}+g_{n},$
where $g_{i}\in\mathbb{F}_{q}[x_{i+1},\ldots,x_{n}]$ and $\deg(g_{i})\leq d_{i}$.
4 The degree of regularity of a random system
In this section, we establish an upper bound for the degree of regularity of an algebraically random system $\mathcal{F}$ consisting of $m$ polynomials of equal degree $D$. In combination with Theorem 3 and Theorem 4, this provides us with an upper bound for the solving degree of a system of algebraically random polynomials of the same degree.
Remark 15.
Notice that $\mathcal{F}\in\mathcal{U}_{\{d,\ldots,d\}}$ for a given $d>0$ if and only if $T\circ\mathcal{F}\circ S\in\mathcal{U}_{\{d,\ldots,d\}}$. In other words, when deciding whether a system of polynomials of equal degree is algebraically random, one can safely ignore the random linear transformations $S$ and $T$ used to disguise the internal system $\mathcal{F}$. This also shows that, for system whose equations are all of the same degree, being an algebraically random system is an intrinsic property of the system and it is not affected by the invertible linear transformations used to disguise the system.
In [10, Section 4], the authors provide an upper bound for the degree of regularity of a system of quadratic polynomials which contains a regular sequence. In this section, we follow the same basic approach and extend it to systems of polynomials of the same degree and systems of polynomials of the same degree to which one adds the field equations. The cases that we treat in this paper are technically more challenging and require the use of more sophisticated results from commutative algebra.
We start by introducing the family of lex-segment ideals. A conjecture by Eisenbud, Green, and Harris will allow us to reduce to these ideals, when estimating the regularity of ideals generated by algebraically random systems. Throughout the section, we fix the lexicographic term order on $R$ with $x_{1}>x_{2}>\ldots>x_{n}$.
@@ 4 The degree of regularity of a random system
For any lex-segment ideal $L$ generated in degree $3$ with $\dim(L_{3})\leq 7$, one has $L\subseteq\mathcal{L}$. Hence one may write
@@
as in Definition 16 and choose e.g. $L=(x^{3},x^{2}y,x^{2}z,xy^{2},xyz,xz^{2})$, or $L=(x^{3})$. However, Notation 17 prescribes that we choose the largest $L$ with respect to containment, i.e.
@@
The next conjecture appears as [19, Conjecture ($V_{m}$)]. It has been settled in several cases and it is widely believed to hold within the commutative algebra community. For an introduction to the conjecture and an excellent survey of known cases, we refer the interested reader to [17]. Here we state it in a weak form, which is what we need in the sequel.
@@
In order to estimate the degree of regularity of our systems, we use the following result by Caviglia and De Stefani.
Proposition 20 ([16, Lemma 2.3]).
@@
Our first result is an explicit bound for the degree of regularity of an algebraically random system of polynomials of the same degree. In order to make the proof more readable, we introduce the following notation.
Notation 21.
@@
In the next theorem, we provide an explicit formula for the degree of regularity of an LPP ideal with given parameters. Thanks to Conjecture 19, this yields an upper bound for the degree of regularity of an algebraically random system of polynomials of the same degree. In the statement of the theorem, $\sigma_{k,t}$ is the position of the smallest $(D,k,D-t)$-type monomial in the ordered list of monomials of degree $D$ different from $x_{1}^{D},\ldots,x_{n}^{D}$, sorted in decreasing lexicographic order. In particular, $\sigma_{1,0}=0$ and $\sigma_{n-1,D-1}$ is the number of monomials of degree $D$ different from $x_{1}^{D},\ldots,x_{n}^{D}$.
@@
Assume that Conjecture 19 holds. Let =$\mathcal{F}=\{f_{1},\ldots,f_{m}\}\subseteq R$ be a polynomial system and assume without loss of generality that $f_{1}^{\operatorname{\mathrm{top}}},\ldots,f_{m}^{\operatorname{\mathrm{top}}}$ are linearly independent of degree $D$. If $m=n$, then let $k=0$, $t=D-1$. Else, let $11\leq k\leq n-$ 1 and $11\leq t\leq D-$ 1 be such that $m-n$ belongs to the interval $(\sigma_{k,t-1},\sigma_{k,t}]$, where
$\sigma_{k,t}=\sum_{i=1}^{k}\sum_{j=1}^{D-1}\binom{n-i-1+j}{j}-\sum_{j=t+1}^{D-1}\binom{n-k-1+j}{j}.$
If $\mathcal{F}$ is an algebraically random polynomial system, then
@@
For $11\leq k\leq n-$ 1 and $10\leq t\leq D-$ 1, the number of $(D,k,D-t)$-type monomials is
$\dim(\mathbb{F}_{q}[x_{k+1},\ldots,x_{n}]_{t})=\binom{n-k-1+t}{t},$
@@
$\begin{split}\sigma_{k,t}&=\sum_{i=1}^{k-1}\sum_{j=1}^{D-1}\binom{n-i-1+j}{j}+\sum_{j=1}^{t}\binom{n-k-1+j}{j}\\ &=\sum_{i=1}^{k-1}\sum_{j=1}^{D-1}\dim(\mathbb{F}_{q}[x_{i+1},\ldots,x_{n}])_{j}+\sum_{j=1}^{t}\dim(\mathbb{F}_{q}[x_{k+1},\ldots,x_{n}])_{j}\end{split}$
is the number of degree $D$ monomials in $\mathbb{F}_{q}[x_{1},\ldots,x_{n}]$ different from $x_{1}^{D},\ldots,x_{n}^{D}$ and bigger than or equal to $x_{k}^{D-t}x_{n}^{t}$, the smallest $(D,k,D-t)$-type monomial. In other words, the monomial $x_{k}^{D-t}x_{n}^{t}$ is in position $\sigma_{k,t}$ in the ordered list of degree $D$ monomials in $\mathbb{F}_{q}[x_{1},\ldots,x_{n}]/(x_{1}^{D},\ldots,x_{n}^{D})$. Notice moreover that $\sigma_{1,0}=0$ and $\sigma_{k,0}=\sigma_{k-1,D-1}$ for $12\leq k\leq n-$ 1.
If $u$ is the smallest monomial in $\mathcal{L}_{D}/\langle x_{1}^{D},\ldots,x_{n}^{D}\rangle$, then $u$ is in position $m-n$ in the ordered list of degree $D$ monomials in $\mathbb{F}_{q}[x_{1},\ldots,x_{n}]/(x_{1}^{D},\ldots,x_{n}^{D})$. If $\sigma_{k,t-1}<m-n\leq\sigma_{k,t}$ for some $11\leq k\leq n-$ 1 and $11\leq t\leq D-$ 1, then $u$ is a $(D,k,D-t)$-type monomial. Since $0=\sigma_{1,0}<m-n\leq\dim(\mathbb{F}_{q}[x_{1},\ldots,x_{n}]_{D})/(x_{1}^{D},\ldots,x_{n}^{D})_{D}=\sigma_{n-1,D-1}$, then $m-n$ always belong to one of the intervals above.
The ideal $L$ is generated by the degree $D$ monomials which are greater than or equal to $u$, unless the monomial following $u$ in lexicographic decreasing order is a pure power. In that case, $u=x_{k}x_{n}^{D-1}$ for some $11\leq k\leq n-$ 1, i.e. $m-n=\sigma_{k,D-1}$, and the smallest degree $D$ monomial in $L$ is $x_{k+1}^{D}$. Then $L$ is generated by the degree $D$ monomials which are greater than or equal to $x_{k+1}^{D}$, which is a $(D,k+1,D)$-type monomial. In both situations
@@
In this example we show how to compute the bound from Theorem 22 for concrete choices of the parameters. Let $n=6$ and $D=3$, so =$\mathcal{F}=\{f_{1},\ldots,f_{m}\}\subseteq\mathbb{F}_{q}[x_{1},\ldots,x_{6}]_{3}$. Then $1\leq k\leq 5$ and $0\leq t\leq 2$. The values of $\sigma_{k,t}$ are
@@
If $m=12$, then $m-n=6$ and $\sigma_{1,1}<6\leq\sigma_{1,2}$. Hence $k=1$, $t=2$, and $d_{\operatorname{reg}}(\mathcal{F})\leq 11.$
If $m=42$, then $m-n=36$ and $\sigma_{3,0}<36\leq\sigma_{3,1}$. Hence $k=3$, $t=1$, and $d_{\operatorname{reg}}(\mathcal{F})\leq 8.$
If $m=54$, then $m-n=48$ and $\sigma_{4,1}<48\leq\sigma_{4,2}$. Hence $k=4$, $t=2$, and $d_{\operatorname{reg}}(\mathcal{F})\leq 5.$
@@
The upper bound of Theorem 22 is decreasing as a function of $m$, as one would expect. In particular, as $m-n$ passes from an interval $(\sigma_{k,t-1},\sigma_{k,t}]$ to the next, the upper bound decreases by one. The largest value of the bound is obtained in the case $m=n$, which corresponds to $\mathcal{F}^{\operatorname{\mathrm{top}}}$ being a regular sequence. In this case, the value for the bound is well-know and is $n(D-1)+1$. The smallest value for the bound is obtained in the case $m-n=\sigma_{n-1,D-1}$, which corresponds to $\langle\mathcal{F}^{\operatorname{\mathrm{top}}}\rangle+\langle x_{1}^{D},\ldots,x_{n}^{D}\rangle=R_{D}$. In this case, the value for the bound is easily seen to be $D$.
Remark 25.
The upper bound produced in Theorem 22 is sharp for all values of $m,n,D$. In fact, it is met by any system $\mathcal{F}$ such that $(f_{1}^{\operatorname{\mathrm{top}}},\ldots,f_{m}^{\operatorname{\mathrm{top}}})$ is a $(D,\ldots,D;D)$-LPP ideal.
Combining Theorem 22 and Theorem 4, we obtain the following.
@@
Let $\mathcal{F}\subseteq R$ be an algebraically random system of degree $D$ polynomials. If $m=n$, then let $k=0$, $t=D-1$. Else, let $11\leq k\leq n-$ 1 and $11\leq t\leq D-$ 1 be such that $m-n$ belongs to the interval $(\sigma_{k,t-1},\sigma_{k,t}]$. If Conjecture 19 holds, then
@@
We now wish to apply Theorem 22 to a system which contains the field equations.
@@
After reducing the equations of $\mathcal{F}$ modulo the field equations, they have degree at most $q-1$ in each variable, hence total degree at most $n(q-1)$. Therefore, when adding the field equations to a polynomial system of degree $D$, we may always assume that
@@
Combining Theorem 22 and Theorem 3, one obtains the following. By Remark 27, we may assume without loss of generality that $D\leq n(q-1)$.
Corollary 28.
Assume that Conjecture 19 holds. Let $\mathcal{F}\subseteq R$ be an algebraically random system of degree $D$ polynomials, with $D\leq n(q-1)$. If $m=n$, then let $k=0$, $t=D-1$. Else, let $11\leq k\leq n-$ 1 and $11\leq t\leq D-$ 1 be such that $m-n$ belongs to the interval $(\sigma_{k,t-1},\sigma_{k,t}]$. If $d_{\operatorname{reg}}(\mathcal{F})\geq q$, then
$\operatorname{\mathrm{solv.deg}}^{s}(\mathcal{F}\cup\{x_{1}^{q}-x_{1},\ldots,x_{n}^{q}-x_{n}\})\leq 2\min\{n(q-1),(n-k+1)(D-1)-t\}.$
Proof.
First, since $(x_{1}^{q},\ldots,x_{n}^{q})_{n(q-1)+1}=R_{n(q-1)+1}$, one has
$d_{\operatorname{reg}}(\mathcal{F}^{\operatorname{\mathrm{top}}}\cup\{x_{1}^{q},\ldots,x_{n}^{q}\})\leq n(q-1)+1.$
If $D<q$, then
$(\mathcal{F}^{\operatorname{\mathrm{top}}}\cup\{x_{1}^{q},\ldots,x_{n}^{q}\})_{q-1}=(\mathcal{F}^{\operatorname{\mathrm{top}}})_{q-1}\neq R_{q-1},$
since $d_{\operatorname{reg}}(\mathcal{F})\geq q$ by assumption. Therefore, $d_{\operatorname{reg}}(\mathcal{F}\cup\{x_{1}^{q}-x_{1},\ldots,x_{n}^{q}-x_{n}\})\geq q=\max\{D,q\}$. If $q\leq D\leq n(q-1)$, then
$(\mathcal{F}^{\operatorname{\mathrm{top}}}\cup\{x_{1}^{q},\ldots,x_{n}^{q}\})_{D-1}=(x_{1}^{q},\ldots,x_{n}^{q})_{D-1}\neq R_{D-1},$
hence $d_{\operatorname{reg}}(\mathcal{F}\cup\{x_{1}^{q}-x_{1},\ldots,x_{n}^{q}-x_{n}\})\geq D=\max\{D,q\}$. This shows that the assumptions of Theorem 3 are satisfied. The thesis now follows by combining Theorem 22, Theorem 3, and the observation that
$d_{\operatorname{reg}}(\mathcal{F}\cup\{x_{1}^{q}-x_{1},\ldots,x_{n}^{q}-x_{n}\})\leq d_{\operatorname{reg}}(\mathcal{F}).$
Remark 29.
While the estimates of Corollary 26 and of Corollary 28 hold for every $D$, they are most relevant for $D\leq q$. In fact, for $D>q$ we obtain tighter upper bounds on the degree of regularity - hence on the solving degree - of $\mathcal{F}\cup\{x_{1}^{q}-x_{1},\ldots,x_{n}^{q}-x_{n}\}$ in Theorem 34 by inspecting the degree $D$ part of the system $\mathcal{F}\cup\{x_{1}^{q}-x_{1},\ldots,x_{n}^{q}-x_{n}\}\in\mathcal{U}_{\{q,\ldots,q\}}$. This corresponds to the fact that, whenever the degree of the equations of the system is larger than or comparable to the field size, it is convenient to add the field equations to the system before computing a Gröbner basis.
Next we estimate the degree of regularity and the solving degree of $\mathcal{F}\cup\{x_{1}^{q}-x_{1},\ldots,x_{n}^{q}-x_{n}\}$ in the case when $D\geq q$. By Remark 27, we may assume that $D\leq n(q-1)$.
@@
Let =$\mathcal{F}=\{f_{1},\ldots,f_{m}\}$ be a polynomial system of degree $q\leq D\leq n(q-1)$. Since $(\mathcal{F}^{\operatorname{\mathrm{top}}}\cup\{x_{1}^{q},\ldots,x_{n}^{q}\})_{D-1}=(x_{1}^{q},\ldots,x_{n}^{q})_{D-1}\neq R_{D-1}$, then
@@
Remark 31.
Assume that $q\leq D\leq n(q-1)$. A standard Hilbert function computation shows that the number of homogeneous polynomials in $n$ variables of degree $D$ which are linearly independent modulo $(x_{1}^{q},\ldots,x_{n}^{q})$ is
@@
It follows that the assumption that $f_{1}^{\operatorname{\mathrm{top}}},\ldots,f_{m}^{\operatorname{\mathrm{top}}}$ are linearly independent modulo $(x_{1}^{q},\ldots,x_{n}^{q})$ holds on a dense open set, whenever
If inequality (2) is not satisfied, then $f_{1}^{\operatorname{\mathrm{top}}},\ldots,f_{m}^{\operatorname{\mathrm{top}}}$ cannot be linearly independent.
The simplest case to treat is that of very overdetermined systems, more specifically the case when $f_{1}^{\operatorname{\mathrm{top}}},\ldots,f_{m}^{\operatorname{\mathrm{top}}}$ are too many to be linearly independent modulo $(x_{1}^{q},\ldots,x_{n}^{q})$. This happens when $m$ is larger than the bound from Remark 31. The next proposition shows that, in such a situation, the degree of regularity is equal to the degree of the equations of the system.
@@
Let =$\mathcal{F}=\{f_{1},\ldots,f_{m}\}$ be a polynomial system of degree $D$, where $q\leq D\leq n(q-1)$. If $m>\sum_{k=0}^{\lfloor\frac{D}{q}\rfloor}(-1)^{k}\binom{n}{k}\binom{n+D-1-kq}{n-1}$, then there is a dense open set $\mathcal{W}$ such that, if $\mathcal{F}\in\mathcal{W}$, then
@@
$m>\sum_{k=0}^{\lfloor\frac{D}{q}\rfloor}(-1)^{k}\binom{n}{k}\binom{n+D-1-kq}{n-1}=\dim(R/(x_{1}^{q},\ldots,x_{n}^{q}))_{D},$

### v3

@@ 4 Bounds on the degree of regularity
Let $\mathcal{F}$ be a system of polynomials with $f_{1}^{\operatorname{\mathrm{top}}},\ldots,f_{m}^{\operatorname{\mathrm{top}}}$ linearly independent of degree $D$. Let $S$ and $T$ be invertible square matrices of size $n$ and $m$. Then $(T\circ\mathcal{F}\circ S)^{\operatorname{\mathrm{top}}}=T\circ\mathcal{F}^{\operatorname{\mathrm{top}}}\circ S$, so $\mathcal{F}$ is regular in degree $D$ if and only if $T\circ\mathcal{F}\circ S$ is regular in degree $D$. In other words, when deciding whether a system of polynomials of equal degree $D$ is regular in degree $D$, one can ignore the random linear transformations $S$ and $T$ used to disguise the internal system $\mathcal{F}$.
In [12, Section 4], the authors provide an upper bound for the degree of regularity of a system of quadratic polynomials that contains a regular sequence. In this section, we follow the same basic approach and extend it to systems of polynomials of the same degree and systems of polynomials of the same degree to which one adds the field equations. The cases we treat in this paper are technically more challenging and require the use of more sophisticated results from commutative algebra.
We start by introducing the family of lex-segment ideals. A conjecture by Eisenbud, Green, and Harris will allow us to reduce to these ideals, when estimating the regularity of ideals generated by systems which are regular in degree $D$, for some $D>0$. Throughout the section, we fix the lexicographic term order on $R$ with $x_{1}>x_{2}>\ldots>x_{n}$.
@@
For any lex-segment ideal $L$ generated in degree $3$ with $\dim(L_{3})\leq 7$, one has $L\subseteq\mathcal{L}$. Hence, one may write
@@
as in Definition 17 and choose, e.g., $L=(x^{3},x^{2}y,x^{2}z,xy^{2},xyz,xz^{2})$, or $L=(x^{3})$. However, Notation 18 prescribes that we choose the largest $L$ with respect to containment, i.e.,
@@
The next conjecture appears as [27, Conjecture ($V_{m}$)]. See also [37, Conjecture 2.6] for an equivalent formulation, that is closer to the formulation that we adopt here. The conjecture was settled in several cases and it is widely believed to hold within the commutative algebra community. For an introduction to the conjecture and an excellent survey of known cases, we refer the interested reader to [21]. Here we state it in a weak form, which is what we need in the sequel.
@@
The conjecture is motivated by the following observation. A classical result by Macaulay shows that, among all homogeneous ideals with the same Hilbert function in a given degree $C$, the lex-segment generated in degree $C$ has the smallest Hilbert function in the next degree. If one restricts to ideals that contain a regular sequence in given degrees, the natural analog of the lex-segment is the LPP ideal from Notation 18. For ideals that contain a regular sequence of maximum cardinality, minimum growth of the Hilbert function corresponds to maximum Castelnuovo-Mumford regularity. In order to compute the degree of regularity of an LPP ideal, we use the following result by Caviglia and De Stefani.
Proposition 21 ([20, Lemma 2.3]).
@@
Our first result is an explicit upper bound for the degree of regularity of a system of polynomials that is regular in degree $D$. In order to make the proof more readable, we introduce the following notation.
Notation 22.
@@
In the next theorem, we provide an explicit upper bound for the degree of regularity of a system that is regular in degree $D$. The result is conditional on Conjecture 20. Notice that the result can be applied to practically any cryptographic system of equations of the same degree. In fact, even if the system of interest is not zero-dimensional, one can assign random values to some of the variables in order to obtain a zero-dimensional system, which is then solved for the remaining variables.
@@
Assume that Conjecture 20 holds. Let $\mathcal{F}=\{f_{1},\ldots,f_{m}\}\subseteq R$ be a polynomial system and assume that $f_{1}^{\operatorname{\mathrm{top}}},\ldots,f_{m}^{\operatorname{\mathrm{top}}}$ are linearly independent of degree $D$. If $m=n$, then let $k=0$, t=D1t=D-1. Else, let $1\leq k\leq n-1$ and $1\leq t\leq D-1$ be such that mnm-n belongs to the interval $(\sigma_{k,t-1},\sigma_{k,t}]$, where
$\sigma_{k,t}=\sum_{i=1}^{k}\sum_{j=1}^{D-1}\binom{n-i-1+j}{j}-\sum_{j=t+1}^{D-1}\binom{n-k-1+j}{j}\;\;\mbox{ for }\;0\leq t\leq D-1.$
If $d_{\operatorname{reg}}(\mathcal{F})<+\infty$, then
@@
In order to apply Proposition 21, we need to determine the value of $k$ and $t_{k}$ for the monomial in position mnm-n in the ordered list of degree $D$ monomials in $\mathbb{F}[x_{1},\ldots,x_{n}]/(x_{1}^{D},\ldots,x_{n}^{D})$. As we will show, these values are the unique $k$ and $t_{k}$ such that mnm-n belongs to the interval $(\sigma_{k,D-t_{k}-1},\sigma_{k,D-t_{k}}]$.
For $1\leq k\leq n-1$ and $0\leq t\leq D-1$, the number of (D,k,Dt)(D,k,D-t)-type monomials is
$\dim(\mathbb{F}[x_{k+1},\ldots,x_{n}]_{t})=\binom{n-k-1+t}{t}$,
@@
$\begin{split}\sigma_{k,t}&=\sum_{i=1}^{k-1}\sum_{j=1}^{D-1}\binom{n-i-1+j}{j}+\sum_{j=1}^{t}\binom{n-k-1+j}{j}\\ &=\sum_{i=1}^{k-1}\sum_{j=1}^{D-1}\dim(\mathbb{F}[x_{i+1},\ldots,x_{n}])_{j}+\sum_{j=1}^{t}\dim(\mathbb{F}[x_{k+1},\ldots,x_{n}])_{j}\end{split}$
is the number of degree $D$ monomials in $\mathbb{F}[x_{1},\ldots,x_{n}]$ different from $x_{1}^{D},\ldots,x_{n}^{D}$ and bigger than or equal to $x_{k}^{D-t}x_{n}^{t}$, the smallest (D,k,Dt)(D,k,D-t)-type monomial. In other words, the monomial $x_{k}^{D-t}x_{n}^{t}$ is in position $\sigma_{k,t}$ in the ordered list of degree $D$ monomials in $\mathbb{F}[x_{1},\ldots,x_{n}]/(x_{1}^{D},\ldots,x_{n}^{D})$. Notice moreover that $\sigma_{1,0}=0$ and $\sigma_{k,0}=\sigma_{k-1,D-1}$ for $2\leq k\leq n-1$.
If $u$ is the smallest monomial in $\mathcal{L}_{D}/\langle x_{1}^{D},\ldots,x_{n}^{D}\rangle$, then $u$ is in position mnm-n in the ordered list of degree $D$ monomials in $\mathbb{F}[x_{1},\ldots,x_{n}]/(x_{1}^{D},\ldots,x_{n}^{D})$. If $\sigma_{k,t-1}<m-n\leq\sigma_{k,t}$ for some $1\leq k\leq n-1$ and $1\leq t\leq D-1$, then $u$ is a (D,k,Dt)(D,k,D-t)-type monomial. Since $0=\sigma_{1,0}<m-n\leq\dim(\mathbb{F}[x_{1},\ldots,x_{n}]_{D})/(x_{1}^{D},\ldots,x_{n}^{D})_{D}=\sigma_{n-1,D-1}$, then mnm-n always belong to one of the intervals above.
The ideal $L$ is generated by the degree $D$ monomials which are greater than or equal to $u$, unless the monomial following $u$ in lexicographic decreasing order is a pure power. In that case, $u=x_{k}x_{n}^{D-1}$ for some $1\leq k\leq n-1$, i.e. $m-n=\sigma_{k,D-1}$, and the smallest degree $D$ monomial in $L$ is $x_{k+1}^{D}$. Then $L$ is generated by the degree $D$ monomials which are greater than or equal to $x_{k+1}^{D}$, which is a $(D,k+1,D)$-type monomial. In both situations
@@
We stress that the upper bound from Theorem 23 is sharp for all values of $m,n,D$. In fact, it is met by any system $\mathcal{F}$ such that $(f_{1}^{\operatorname{\mathrm{top}}},\ldots,f_{m}^{\operatorname{\mathrm{top}}})$ is a $(D,\ldots,D;D)$-LPP ideal.
In the next example we show how to compute the bound from Theorem 23 for concrete choices of the parameters.
@@
Let $n=6$ and $D=3$, so $\mathcal{F}=\{f_{1},\ldots,f_{m}\}\subseteq\mathbb{F}_{q}[x_{1},\ldots,x_{6}]_{3}$. Then $1\leq k\leq 5$ and $0\leq t\leq 2$. The values of $\sigma_{k,t}$ are
@@
If $m=12$, then mn=6m-n=6 and $\sigma_{1,1}<6\leq\sigma_{1,2}$. Hence $k=1$, $t=2$, and $d_{\operatorname{reg}}(\mathcal{F})\leq 11.$
If $m=42$, then mn=36m-n=36 and $\sigma_{3,0}<36\leq\sigma_{3,1}$. Hence $k=3$, $t=1$, and $d_{\operatorname{reg}}(\mathcal{F})\leq 8.$
If $m=54$, then mn=48m-n=48 and $\sigma_{4,1}<48\leq\sigma_{4,2}$. Hence $k=4$, $t=2$, and $d_{\operatorname{reg}}(\mathcal{F})\leq 5.$
@@
The upper bound of Theorem 23 is decreasing as a function of $m$, as one would expect. In particular, as mnm-n passes from an interval $(\sigma_{k,t-1},\sigma_{k,t}]$ to the next, the upper bound decreases by one. The largest value of the bound is obtained in the case $m=n$, which corresponds to $\mathcal{F}^{\operatorname{\mathrm{top}}}$ being a regular sequence. In this case, the value for the bound is well-known and is n(D1)+1n(D-1)+1. The smallest value for the bound is obtained in the case $m-n=\sigma_{n-1,D-1}$, which corresponds to $\langle\mathcal{F}^{\operatorname{\mathrm{top}}}\rangle+\langle x_{1}^{D},\ldots,x_{n}^{D}\rangle=R_{D}$. In this case, the value of the bound is $D$.
In order to provide a concrete example, we apply Theorem 23 to the Cubic Simple Matrix encryption scheme proposed in [26]. This provides the only proven bound on the degree of regularity of this system, to the best of our knowledge.
Example 26.
The Cubic Simple Matrix encryption scheme consists of $m=2n$ cubic equations in $n$ variables, with $n$ a square. To compute the bound of Theorem 23 for this system, we need to find the unique $1\leq k\leq n-1$ and $t\in\{1,2\}$ such that $m-n=n\in(\sigma_{k,t-1},\sigma_{k,t}]$. Since $\sigma_{1,1}=n-1$ and $\sigma_{1,2}=n-1+{n\choose 2}$, we find $k=1$ and $t=2$. Hence, the Cubic Simple Matrix encryption scheme has $d_{\operatorname{reg}}\leq 2n-1$.
We are not aware of any other proven bound on the degree of regularity of this system. However, [26] contains a heuristic analysis for $n\gg 0$ and experimental results for small $n$. In [26, Section 4.2] the authors perform a heuristic analysis and conclude that, for $n$ sufficiently large, the complexity of solving the cubic system using the XL Algorithm should be at least that of solving a random quadratic system of $n$ equations in $n$ variables. The degree of regularity of the latter is $n+1$, provided that the top degree part of the equations is a regular sequence. In the same section, they carry on numerical experiments for $n=4,9,16$, from which they find degree of regularity $5,6,7$, respectively.
Combining Theorem 23 and Theorem 4, we obtain the following. A related result concerning the solving degree of a standard algorithm will be proven after Theorem 33.
@@
Let $\mathcal{F}=\{f_{1},\ldots,f_{m}\}\subseteq R$ be a system of polynomials such that $f_{1}^{\operatorname{\mathrm{top}}},\ldots,f_{m}^{\operatorname{\mathrm{top}}}$ are linearly independent of degree $D$. Assume that $d_{\operatorname{reg}}(\mathcal{F})<+\infty$. If $m=n$, then let $k=0$, t=D1t=D-1. Else, let $1\leq k\leq n-1$ and $1\leq t\leq D-1$ be such that mnm-n belongs to the interval $(\sigma_{k,t-1},\sigma_{k,t}]$. If Conjecture 20 holds, then
@@
While the estimates of Theorem 23 and Corollary 27 hold for every $D$, they are most relevant for $D<q$. In fact, if $D\geq q$, the bound on the solving degree of a mutant system obtained using Theorem 4 does not increase if one adds the field equations to the system. Moreover, one always has $d_{\operatorname{reg}}(\mathcal{F}\cup\{x_{1}^{q}-x_{1},\ldots,x_{n}^{q}-x_{n}\})\leq d_{\operatorname{reg}}(\mathcal{F})$. Therefore, if $D\geq q$, one should add the field equations to the system before solving it.
We now bound the degree of regularity and the solving degree of a system of the form $\mathcal{F}\cup\{x_{1}^{q}-x_{1},\ldots,x_{n}^{q}-x_{n}\}$, where $\mathcal{F}$ is a system of $m$ equations of degree $D$ in $n$ variables. This is especially relevant in the case $\mathbb{F}=\mathbb{F}_{q}$, but the results that we will prove hold (and are stated) over an arbitrary field $\mathbb{F}$. Because of Remark 28, we focus on the case when $D\geq q$. After reducing the equations of $\mathcal{F}$ modulo the field equations, they have degree at most q1q-1 in each variable, hence total degree at most n(q1)n(q-1). Therefore, in the sequel we may assume without loss of generality that
@@
Let $\mathcal{F}=\{f_{1},\ldots,f_{m}\}$ be a polynomial system of degree $D\leq n(q-1)$. Since $(\mathcal{F}^{\operatorname{\mathrm{top}}}\cup\{x_{1}^{q},\ldots,x_{n}^{q}\})_{D-1}=(x_{1}^{q},\ldots,x_{n}^{q})_{D-1}\neq R_{D-1}$, then
@@
Lemma 30.
Let $\mathcal{F}=\{f_{1},\ldots,f_{m}\}\subseteq R$ with $\deg(f_{1})=\ldots=\deg(f_{m})=D$. Assume that $q\leq D\leq n(q-1)$. If $f_{1}^{\operatorname{\mathrm{top}}},\ldots,f_{m}^{\operatorname{\mathrm{top}}}$ are linearly independent of degree $D$, then
Conversely, if inequality (3) is satisfied, then the set of systems $\mathcal{F}$ such that that $f_{1}^{\operatorname{\mathrm{top}}},\ldots,f_{m}^{\operatorname{\mathrm{top}}}$ are linearly independent modulo $(x_{1}^{q},\ldots,x_{n}^{q})$ is open and nonempty.
Proof.
A standard Hilbert function computation shows that the number of homogeneous polynomials in $n$ variables of degree $D$ which are linearly independent modulo $(x_{1}^{q},\ldots,x_{n}^{q})$ is
@@
Therefore, if $f_{1}^{\operatorname{\mathrm{top}}},\ldots,f_{m}^{\operatorname{\mathrm{top}}}$ are linearly independent, then inequality (3) is satisfied. The set of systems $\mathcal{F}$ as in the statement is always open, since linear dependence is a rank condition, hence a closed condition. Suppose then that inequality (3) is satisfied. Then there exist $g_{1},\ldots,g_{m}\in(R/(x_{1}^{q},\ldots,x_{n}^{q}))_{D}$ linearly independent, so the set of systems $\mathcal{F}$ in the statement is nonempty. ∎
The simplest case is that of very overdetermined systems, more specifically the case when $f_{1}^{\operatorname{\mathrm{top}}},\ldots,f_{m}^{\operatorname{\mathrm{top}}}$ are too many to be linearly independent modulo $(x_{1}^{q},\ldots,x_{n}^{q})$. This happens when $m$ is larger than the bound from Lemma 30. It is easy to show that, in such a situation, the degree of regularity is equal to the degree of the equations of the system.
@@
Let $\mathcal{F}=\{f_{1},\ldots,f_{m}\}$ be a polynomial system of degree $D$, where $q\leq D\leq n(q-1)$. If $m\geq\sum_{k=0}^{\lfloor\frac{D}{q}\rfloor}(-1)^{k}\binom{n}{k}\binom{n+D-1-kq}{n-1}$, then there is a dense open set $\mathcal{W}$ such that, if $\mathcal{F}\in\mathcal{W}$, then
@@
$m\geq\sum_{k=0}^{\lfloor\frac{D}{q}\rfloor}(-1)^{k}\binom{n}{k}\binom{n+D-1-kq}{n-1}=\dim(R/(x_{1}^{q},\ldots,x_{n}^{q}))_{D},$